OpenStudy (anonymous):

Please help!! A box contains 6 nuts, 8 bolts, and 4 screws. If 3 objects are selected in succession randomly, what is the probability of selecting a nut, then a bolt, then a screw, if replacement occurs each time? A. 8/243 B. 16/243 C. 1/27 D. 1

jimthompson5910 (jim_thompson5910):

Hint: Assuming the selections are independent, you can say P(Nut, bolt, screw) = P(Nut)*P(bolt)*P(screw)

jimthompson5910 (jim_thompson5910):

let me know if this helps or not

OpenStudy (anonymous):

so it would be P(6/17) * P(8/17) * P(4/17)?

jimthompson5910 (jim_thompson5910):

lose the P's and turn all the 17's into 18 There are 6+8+4 = 18 items total (and not 17)

jimthompson5910 (jim_thompson5910):

so you would have (6/18) * (8/18) * (4/18) as your next step

OpenStudy (anonymous):

Okay and then you get 192/5832

jimthompson5910 (jim_thompson5910):

now reduce as much as possible

OpenStudy (anonymous):

3/729?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

oh wait it would be 3/3 = 1

jimthompson5910 (jim_thompson5910):

I would divide by 2 first 192/2 = 96 5832/2 = 2916 So 192/5832 = 96/2916 we can go further

jimthompson5910 (jim_thompson5910):

96/2 = 48 2916/2 = 1458 So 96/2916 = 48/1458

jimthompson5910 (jim_thompson5910):

Since 2 works a bunch of times, let's try 4 48/4 = 12 1458/4 = 364.5 so 4 doesn't work, so let's try 2 again ------------------------------------------------------- Anyways, you do this trial and error a bunch of times The best way to do it really is to find the GCF of 192 and 5832, which is 24, and divide each term by that 192/24 = 8 5832/24 = 243

jimthompson5910 (jim_thompson5910):

So in the end, 192/5832 reduces to 8/243

OpenStudy (anonymous):

Oh okay I get it now. Thank you

jimthompson5910 (jim_thompson5910):

you're welcome